OpenStudy (anonymous):

You have already invested $450 in a stock with an annual return of 11%. How much of an additional $900 should be invested at 12% and how much at 6% so that the total return on the entire $1,350 is 9%? (Round each answer to the nearest cent.)

OpenStudy (anonymous):

9% on 1350 = 121.50 11% on 450 = 49.50 SUBTRACT 72.00 $ Let the amount invested at 12% be x $ Then the amount invested at 6% will be 900-x $ 0.12x + 0.06( 900 -x) = 72 0.12x + 54 -0.06x = 72 0.06x = 18 x = 18/0.06 = 300 $ ANSWER 300$ at 12% and 600$ at 6% CHECK 300 X 12/100 - = 36 600 X 6/100 = 36 450 X 11/100= 49.50 ADD 121.50 $ 121.50 /1350 X 100 = 9%

OpenStudy (anonymous):

Your amazing! Thank you!

OpenStudy (anonymous):

np.